### 10 Jan 2024 v1

# ANGO: A Next-Level Evaluation Benchmark For Generation-Oriented Language Models In Chinese Domain

Abstract
Recently, various Large Language Models (LLMs) evaluation datasets have emerged, but most of them have issues with distorted rankings and difficulty in model capabilities analysis. Addressing these concerns, this paper introduces ANGO, a Chinese multi-choice question evaluation benchmark. ANGO proposes Keypoint categorization standard for the first time, each question in ANGO can correspond to multiple keypoints, effectively enhancing interpretability of evaluation results. Base on performance of real humans, we build a quantifiable question difficulty standard and divide ANGO questions into 9 difficulty levels, which provide more precise guidance for model training. To minimize data leakage impact and fully leverage ANGO’s innovative features, we have engineered exclusive sampling strategies and a new evaluation framework that support swift testset iteration. Our experiments demonstrate that ANGO poses a stronger challenge to models and reveals more details in evaluation result compared to existing benchmarks.
1 Introduction
The field of Natural Language Processing (NLP) has witnessed significant advancements in recent years, largely attributed to the development of the LLM which possesses a vast number of parameters and trains on extensive corpus. A pivotal breakthrough is the introduction of the Transformer architecture, which incorporates the Attention mechanism [1]. This architecture facilitates the modeling of long-range dependencies in textual data and enables parallel training. Subsequently, pioneering models such as BERT [2] emerged, setting new standards for self-supervised pretraining. More recently, decoder-only models like GPT [3] have gained widespread adoption as the foundation for constructing LLMs.
Figure 1: ANGO: A Next-level evaluation benchmark for Generation-Oriented language models
As LLMs have progressed, the benchmarks used to evaluate them have also evolved. Initial works primarily focused on assessing Natural Language Understanding (NLU) tasks, exemplified by benchmarks like GLUE [4] and its multi-task variant superGLUE [5]. However, as research has delved deeper into the linguistic capabilities of LLMs, specialized benchmarks with multiple-choice question formats have emerged to evaluate specific abilities such as reasoning and inference. Notable examples include MMLU [6], TruthfulQA [7], Big-Bench [8] in English, and AGI-Eval [9], C-Eval [10], CMMLU [11] for the Chinese domain, which are now widely adopted and prominently featured on leaderboards.
Challenges:
Existing multi-choice question benchmarks primarily rely on collecting exam questions to construct datasets and establish categories and difficulty levels based on the Subject covered in these exams. However, this simplistic categorization suffers from the following general problems:
  • Single-Question Single-Subject: Subject-based categories disregard that questions often require proficiency in multiple subjects and overlooking the existence of further subdivisions within each subject.
  • Immeasurable Difficulty: Difficulty levels typically determined by the educational grade associated with the subject, despite significant variations in difficulty for different questions in same subject.
  • Hard To Update: Benchmark data leaks always result in serious distortion[12]. Most of them are challenging to conduct periodic updates due to potential sampling biases.
Because of the aforementioned issues, traditional benchmarks struggle to provide reliable measurements for model’s ability to handle a diverse range of disciplines and fail to accurately assess the performance for complex questions.
Contributions:
We introduce ANGO as a novel Chinese benchmark for evaluation with the format of multiple-choice questions. ANGO apply Keypoint instead of Subject as the criteria to address previous challenges by following improvements:
  • Single-Question Multi-Keypoints: Question in ANGO can correspond to 1-6 different keypoints, and include 988 keypoint combinations. The number of keypoints reach up to 171 distinct keypoints.
  • Quantifiable Difficulty: For difficulty assessment, we introduce a quantifiable function that incorporates real human performance, and divided questions in ANGO into 9 grades.
  • Easy To Update: We design specific strategies for sampling, which meet the balanced keypoints distribution and ensure minimal information loss when reducing the number of few-shot examples.
With the innovative features of ANGO, we build an evaluation framework that can provide more comprehensive and accurate multi-level, multi-perspective performance results for participating models. Furthermore, this framework allows dynamic updates of the testset to ensure the objectivity and authenticity.
2 Data
2.1 Data Collection
Data Source
Our research data is exclusively sourced from the Administrative Proficiency Test (AAT) used in the Chinese civil service examination, which encompasses 34 provinces in China, including both official and mock exams conducted between 2008 and 2023.
The AAT is a comprehensive assessment tool comprised entirely of multiple-choice questions, designed to evaluate the abilities and skills required for practical administrative work. The language model exhibits remarkable performance in generating and comprehending text, and the AAT provides specific and intricate contexts that simulate communication and decision-making processes in the real world. By utilizing the language model to answer questions from the AAT, we can assess its understanding of complex problems, its ability for judgment and reasoning, as well as the accuracy and fluency of its language expression. Furthermore, the AAT covers a wide range of question domains and scenarios, this diversity contributes to evaluating the language model’s language processing capabilities in different domains.
Figure 2: The Data Preprocess Flow: From 674,675 source data to 71,149 clean data
Data Preprocessing
After collecting without any manipulation, we obtained a total of 674,675 records. In order to enhance the quality of our data, we employed a series of simple but efficient preprocess functions, as shown in Figure 2: (1) Mock exams often include previous exam questions, which lead to numerous duplicates. We remove duplicates based on the content of records; (2) Since our primary focus was on NLP instead of Multi-Modal, we removed all records containing pictures in both the questions and options; (3) By manual recognition aided by OCR model, we extracted 5,574 unique formulas from a pool of 34,062 formula images.
2.2 Data Distribution
After completing Data Collection and Data Preprocessing, we divided the records into different keypoints and difficulty levels based on the question content and the domains involved in them.
2.2.1 Keypoint Tree
Figure 3: Keypoint Tree: The portion of the area represents the data size, including 171 keypoints and 4 levels
The keypoint categorization can comprehensively represent the distribution of capabilities required to complete a question. 5 major categories were identified in ANGO as the first-level keypoints.
  • Comprehension&Expression is the core keypoint that permeates the entire process as a fundamental ability.
  • Quantitative Relations showcases computational abilities, thus reflecting the model’s mathematical proficiency.
  • Material Analysis represents the model’s capacity to scrutinize materials, assessing the aptitude for context analysis.
  • Judgement&Inference is designed for logical structure or discernible patterns, demanding inferential ability.
  • General Knowledge serves as a direct manifestation of the model’s knowledge breadth and guarantees reliability.
Considering the inclusion and similarity relations among keypoints, we also constructed a Keypoint Tree consisting of 4 levels and a total of 171 nodes as shown in Figure 3.
2.2.2 Difficulty Measurement
Figure 4: Difficulty Distribution: x-axis stands for difficulty level and y-axis means the question count
The utilization of standardized and objective methods is crucial in measuring the difficulty of questions. Drawing upon the actual scores derived from the AAT and statistical indicators of human performance, ANGO proposes a difficulty score standard:
$S(q)=\alpha S_{q}-\beta Acc_{q}-\gamma\sum_{K}^{K_{s}}Acc_{K}$ (1)
Where $S_{q}$ represents the score assigned to question $q$ in the actual exam, $Acc_{q}$ denotes the human accuracy of question $q$, $K_{s}$ refers to the keypoints combination of question $q$, $K$ represents the keypoint included in $K_{s}$, and $Acc_{K}$ represents the average accuracy of all questions related to $K$. Besides, the $\alpha$, $\beta$ and $\gamma$ are penalty terms.
With this standard, we categorized the difficulty scores into nine distinct levels, as shown in Figure 4, ranging from Level 1 (easiest) to Level 9 (most challenging).
3 Sample Strategy
Prior to commencing, it is essential to establish certain definitions to circumvent any potential misunderstandings:
  • $K_{s}$: Refers to the combination of keypoints present in each record.
  • $K$: Each $K_{s}$ encompasses one or more keypoints, denoted as $K$.
  • $P(K)$: Node path from root node to target node $K$, contains a series of parent nodes $K^{p}$.
  • $G(K)$: Denotes the group of the sub nodes $K^{g}$ which have same parent node with $K$ in the keypoint tree.
3.1 Few-Shot History Sample
Due to the distinct category hierarchy in our keypoint tree architecture and the insufficient number of records for some $K$, we design a dynamic length example sampling strategy that ensures minimal information loss and divides into below 4 steps. The next step will only be taken if there is not any available record in current step.
  • Step 1-$K_{s}$: Assign records as examples with identical $K_{s}$ as examples directly.
  • Step 2-$K$: Iterate through the $K$ within $K_{s}$ as a loop and sample one record associated with each $K$.
  • Step 3-$K^{p}$: Iterate through the parent nodes $K^{p}$ within $P(K)$ and repeat Step 2 on each $K^{p}$.
  • Step 4-$K^{g}$: Reiterate through the parent nodes $K^{p}$ within $P(K)$, repeat Step 2 on each $K^{g}$ of $G(K^{p})$.
Benefits from the multi-step sampling approach, it ensures that the examples in the beginning are more closely aligned with the test record compared to those at the end.
3.2 Testset Sample
Since $K_{s}$ may contain more than one $K$ and the substantial overlap across different $K_{s}$, sampling records for each $K$ or $K_{s}$ will definitely lead to a bad result. To achieve a more balanced distribution, we propose a sample strategy in following steps:
Build $K_{s}$ Vectors:
Represent each $K_{s}$ as a vector with 171-dimensions, denoted as $V_{K}$. Each dimension in this vector corresponds to a specific $K$, $1$ denotes the presence of $K$ in $K_{s}$, while 00 means absence.
Initialize Testset:
After constructing the vectors, randomly select one record for each $K_{s}$ to populate the testset. Subsequently, sum the 988 $V_{K}$ to form a single vector, denoted as $V_{C}$, where each dimension represents the count of the $K$ in the current testset.
Insert and Update:
Firstly, define the number of records that need to be sampled for each $K$ as $N$. Then, measure the distances between $V_{C}$ and all $V_{K}$ with following distance function:
$D(V_{C},V_{K})=\sum log_{N}V_{C}*\frac{V_{K}}{\|V_{K}\|_{1}}$ (2)
Here, ${\|V_{K}\|_{1}}$ is employed to normalize the weight of $K_{s}$, and the logarithmic transformation $log_{N}$ is applied to prevent the inhibition of some $K$ that may only be associated with $K_{s}$ encompassing other high-frequency $K$.
After calculation, select the $K_{s}$ with the minimum distance, randomly choose a record, and insert it into the test. Concurrently, update $V_{C}$ by adding the selected $V_{K}$. Just repeat this step until each $K$ has $N$ records or there are no more records available for $K$.
4 Evaluation
4.1 Metric
We choose accuracy to be the ultimate metric for assessing the performance of our models.
For the base model(Pre-Trained stage), as there are questions with multiple correct options in our dataset, it’s not possible for us to calculate the logits of different options to get the answer, So we simply split the generated output of the model by newline delimiters and extracted the first element as the answer to match the correct answer.
For chat models(Supervised Fine Tune stage), models often failed to provide answers in our expected format. To address this issue, we split the generation output based on Chinese and English tags and designed some regular expressions to capture options with a special format.
Valuable Failure:
Evaluating models solely based on their ability to provide correct answers does not capture the full spectrum of human-like behavior and thinking patterns. While accuracy is important, overlooking the value in identifying common human errors may hinder progress toward building truly human-centric AI. This motivated us to propose new evaluation metrics to address this limitation.
Human Hit means the option predicted by the model is closest to the majority of human results. When the model identifies an option as "Most Wrong", which is the option that humans are most likely to choose incorrectly, and the human accuracy is below 50%, it’s a "Wrong Hit". Conversely, if the model gets the correct answer and the human accuracy is above 50%, referred to as a "Correct Hit". Human Hit is produced by the sum of these two types of Hit and serves as an approximate method to assess whether the model exhibits a tendency similar to the widespread human misconceptions found in the data.
Human Value is obtained by iterating through all the questions in Human Hit using the following formula.
$V(q)=\sum_{i}^{n}\frac{1}{N}\left\{\begin{array}[]{ll}Acc_{q}&Acc_{q}\geq 0.5,\\ 1-Acc_{q}&Acc_{q}<0.5.\end{array}\right.$ (3)
where $Acc_{q}$ represents Human Accuracy of the question, and $n$ denotes the count of Human Hit and $N$ stands for total count.
While accuracy remains paramount, recognizing value in failure opens new possibilities for the model[13]. Human Hit and Human Value aim to foster models that learn and think in a manner increasingly aligned with humanity’s shared imperfections. This anthropic evaluation approach holds promise for developing AI with new features, like Generative Agents[14], better attuned to human needs, values, and experiences.
4.2 Prompt
Similar to the settings of other benchmarks, ANGO adopts the few-shot as an evaluation strategy with history examples sampled in Section 3.1. By exposing the model to a limited number of examples during the evaluation process, we can effectively gauge its generalization and adaptability to novel scenarios. Furthermore, the few-shot evaluation strategy offers practical advantages when it comes to extracting the model’s output options, an example of a prompt is shown in Figure 5.
Figure 5: Prompt Example: English translations are shown below the corresponding Chinese text for better readability.
In particular, we have made a deliberate decision to minimize the inclusion of Chain Of Thought(CoT)[15] in our evaluation procedures, with the exception of the API over which we lack control. This decision stems primarily from prior research[10] that has unequivocally demonstrated the detrimental impact of CoT on model performance.
In cases where the length exceeds this limit, we systematically reduce the number of examples from the end to the beginning. Thanks to our example sampling strategy(Section 3), the examples at the head provide more informative content compared to those at the tail.
4.3 Framework
We devised a range of solutions to optimize the evaluation framework to address the distorting impact of data leaks and malicious intrusions on the testset, which aim to provide more authoritative and trustworthy guidance to the participating models.
  • Seasonal Dynamic Evaluation: Thanks to our optimized sampling strategy, we can periodically refresh a portion of the testset and update the leaderboard so that each season has a distinct testset to prevent data leaks.
  • Confusion of Options Order: To prevent models from relying on spurious patterns in the order of multiple-choice options, we shuffle the order of options for each question before making predictions.
  • Question Elimination: By calculating questions’ accuracy across models and seasons. Questions exceeding 100% accuracy are temporarily removed to ensure sufficient discriminability.
Together, these mechanisms enhance fairness and effectiveness compared to current leaderboards. The seasonal performance history also provides insights into model optimization trends, serving as a guide for further training.
5 Experiment
5.1 Objectives
  • Apply the keypoint categories of ANGO in practice, provide detailed evaluation result, and analyze the model’s performance under different keypoints.
  • Verify the difficulty grading of ANGO, provide a comparison of different grades, and demonstrate the effectiveness of difficulty classification.
  • Test the evaluation framework of ANGO, provide an overall performance of current models, and validate the rationality of the Human Value metric.
5.2 Models
In light of limited computational resources, we were compelled to utilize a single A100 unit for conducting our experiments. As a result, we were obliged to select the following models of relatively smaller scale or models accessible through API services.
GPT: OpenAI’s GPT has long been regarded as a benchmark, which has achieved SOTA in multiple domains and provides services through API. We select GPT-4[16] and ChatGPT[17] as candidates.
LLaMA & LLaMA Variant: LLaMA, which is released by Meta as an open-source LLM, allows researchers to fine-tune models for specific domains. It is the most widely adopted open-source model now. For our evaluation, we choose the LLaMA[18], LLaMA2[19] and some variant models that re-train on the LLaMA base model, such as ChineseLLaMA/ChineseLLaMA2[20], Atom[21].
Chinese-Oriented: In addition to the English-oriented LLMs mentioned above, we have also selected several LLMs that specialize in the Chinese domain. Including Baichuan/Baichuan2[22], ChatGLM/ChatGLM2[23], Qwen[24], InternLM[25], and MOSS[26].
5.3 Results
The evaluation results of ANGO are actually a multidimensional representation, and we have selected three coarse-grained distributions to showcase from it.
5.3.1 Keypoint Level
Figure 6: Performance on level 1 Keypoints in ANGO(dot line is the human performance))
Model All Expression& Material Quantitive JudgementGeneral Comprehension Analysis Relations Inference Knowledge Total 3863 2296 496 485 350 226 Human 62.54 66.82 69.07 40.06 67.08 46.34 GPT-4 60.67 69.5 45.36 36.29 59.43 58.85 Baichuan2-13B 50.52 60.31 29.64 29.28 41.71 55.75 Qwen-7B 49.07 58.74 29.03 26.19 42.57 53.54 Baichuan2-7B 46.22 54.38 30.24 28.25 38.29 50.0 InternLM-7B 45.7 55.16 29.64 28.04 35.14 39.82 Baichuan-13B 44.98 52.55 30.65 30.72 34.86 46.9 ChatGPT 44.02 48.06 38.1 31.55 41.14 47.79 ChatGLM2-6B 39.0 43.57 31.45 28.04 34.86 40.27 Baichuan-7B 38.01 43.92 30.85 26.6 30.29 30.09 ChineseLLaMA2-13B 35.19 38.21 26.61 29.9 33.14 38.05 Atom-7B 34.41 38.21 27.02 27.01 32.86 30.53 ChatGLM-6B 31.75 39.48 17.34 12.78 30.0 29.2 LLaMA2-13B 30.42 33.73 22.58 26.39 30.86 22.12 ChineseLLaMA2-7B 29.98 30.02 31.45 29.69 30.86 26.11 Moss-7B 28.85 30.68 24.6 25.57 28.57 26.11 ChineseLLaMA-13B 27.86 29.89 25.0 24.95 26.86 21.24 LLaMA2-7B 27.55 28.28 27.22 27.01 27.43 22.12 LLaMA-13B 25.43 24.05 26.81 27.63 33.71 18.14 LLaMA-7B 25.01 26.8 23.59 21.24 24.0 19.91 Table 1: Result On Keypoint Tree Level-1: Because a single question may contain more than one keypoint, the total number of keypoint count is higher than the question count
Figure 6: Performance on level 1 Keypoints in ANGO(dot line is the human performance))
Figure 6 clearly reveals that models are able to achieve higher accuracy levels in the Expression&Comprehension and General Knowledge. Because the question formats and content in these two categories tend to be more inclined towards simple content completion, which is relatively well-suited to the pretraining process of language models[3].
In the Judgement&Inference, model performances closely align with the average accuracy. Dive deeper, we can see some conventional logical problems are well-handled by the models[27]. However, there is still a subset of questions in this part that involve pattern-based problems, where models often struggle due to their inability to comprehend the problem’s context.
Mathematics is widely recognized as a challenging task for LLM[28]. This challenge persists as all models demonstrate significantly lower accuracy levels in the domain of Quantitative Relations. In contrast to existing benchmarks, the format of mathematical questions in ANGO tends to rely more heavily on extensive linguistic content, placing a greater emphasis on the combined language and computational capabilities of the models.
Material Analysis is a crucial keypoint of ANGO, characterized by the longer context resulting from the combination of material and question. Due to the extensive context, models often face challenges of attention dispersion, leading to hallucination[29]. Furthermore, this category always involves different abilities, such as analysis, reasoning, and computation. Therefore, the accuracy in this section is significantly much lower than in humans.
Why ChatGPT fail?
Regarding the previously mentioned underperformance of ChatGPT, Table 1 and Figure 6 demonstrates that ChatGPT can achieve better performance in challenging categories such as Material Analysis and Quantitative Relations by leveraging specialized tools and enhanced capabilities.
In terms of Expression&Comprehension and General Knowledge, which represent a significant proportion. Models specifically trained for Chinese achieve better performance level than others. This status can be attributed to differences in the distribution of language features and locally specific knowledge, causing ChatGPT to struggle.
Furthermore, we observed that when faced with challenging questions containing additional information embedded within specific structures and sentence patterns, ChatGPT tends to generate insufficient conditions or unable-to-answer responses rather than accurately selecting the most relevant answer option, which is primarily due to the overreliance on RLHF, [30].
5.3.2 Difficulty Level
Figure 7: Accuracy on different different difficulty levels(dot line is the human performance)
Difficult All Level 7 Level 6 Level 5 Level 4 Level 3 Level 2 Total 1768 236 306 561 470 193 2 Human 62.54 27.59 49.40 60.4 72.76 88.92 96.78 GPT-4 57.69 34.32 47.06 56.68 67.87 80.83 100.0 Baichuan2-13B 47.51 22.46 36.6 48.84 57.66 66.84 50.0 Qwen-7B 46.38 26.27 39.87 49.2 50.64 62.18 100.0 Baichuan2-7B 43.27 24.15 35.62 43.32 51.7 57.51 100.0 ChatGPT 43.21 23.73 33.66 44.92 49.36 61.66 100.0 InternLM-7B 42.36 20.76 36.93 40.29 51.28 61.66 50.0 Baichuan-13B 41.97 24.15 38.56 41.35 49.79 51.3 100.0 ChatGLM2-6B 38.01 22.03 35.29 40.29 41.91 45.6 50.0 Baichuan-7B 35.35 17.8 33.99 34.58 40.43 48.19 100.0 ChineseLLaMA2-13B 35.07 21.61 34.97 34.76 40.64 38.34 100.0 Atom-7B 33.48 20.76 30.07 32.98 39.15 41.97 50.0 ChatGLM-6B 30.32 16.53 24.84 30.66 35.96 40.93 50.0 ChineseLLaMA2-7B 30.2 18.22 30.72 29.95 34.68 33.68 50.0 LLaMA2-13B 30.2 17.37 27.12 29.41 37.02 36.27 50.0 Moss-7B 28.73 16.1 29.74 28.88 31.91 34.72 0.0 ChineseLLaMA-13B 27.71 19.49 27.45 25.67 32.13 33.16 50.0 LLaMA2-7B 27.43 18.64 28.1 27.09 30.64 30.05 50.0 LLaMA-13B 27.04 22.03 27.45 26.56 28.3 30.57 50.0 LLaMA-7B 24.66 13.98 24.84 24.24 28.94 28.5 0.0 Table 2: Result on Different Difficulty
Figure 7: Accuracy on different different difficulty levels(dot line is the human performance)
As evident from the Figure 7, under the difficulty standards set by ANGO, the performance of the models indeed exhibits a strong negative correlation, indicating that higher difficulty levels are associated with lower accuracy. Furthermore, the model that performs well on average does not necessarily perform well across all difficulty levels, but the overall ranking distribution still aligns with the relationship to the average performance, demonstrating the differentiation across difficulty levels in ANGO. All of these findings confirm that the difficulty standard in ANGO is a reasonable and effective quantitative measure and provides a systematic and objective approach to assessing question difficulty in evaluation.
Besides, we also found that models usually work well on high difficulty level questions compared to real human.
Based on the analysis of keypoint performance, it is clear that the LLM’s robust "memory" capabilities equip it to effectively address challenging questions within specific domain of knowledge. In contrast, real human generally possess a restricted comprehension across diverse subject areas, yet they excel in solution with limited information. As a result, the model’s superior performance in high difficulty level questions, which belong to Expression&Comprehension and General Knowledge, and this empowers the model to outperform humans in tackling intricate problems.
5.3.3 Question Level
Figure 8: Acc&Human Value: Dot line is the human performance, ignore the Human Hit of Human(because it’s 100%)
Model Acc % Human Human Human Correct Total Value % Hit % Hit Human 60.5594 70.797 100 1768 1208 1768 GPT-4 57.6923 40.2748 54.3552 961 1020 1768 Baichuan2-13B 47.5113 36.1515 49.2081 870 840 1768 Qwen-7B 46.3801 33.6215 45.871 811 820 1768 ChatGPT 43.6086 33.0455 45.0792 797 771 1768 Baichuan2-7B 43.2692 32.6535 44.5701 788 765 1768 InternLM-7B 42.3643 32.1553 43.7783 774 749 1768 Baichuan-13B 41.9683 31.9635 43.948 777 742 1768 ChatGLM2-6B 38.009 28.4656 39.5928 700 672 1768 Baichuan-7B 35.3507 28.2665 38.9706 689 625 1768 ChineseLLaMA2-13B 35.0679 27.3031 37.7262 667 620 1768 Atom-7B 33.4842 26.0837 35.8032 633 592 1768 ChineseLLaMA2-7B 30.2036 23.3305 32.2964 571 534 1768 LLaMA2-13B 30.2036 23.5473 32.5226 575 534 1768 ChatGLM-6B 29.4118 23.4395 31.8439 563 520 1768 Moss-7B 28.733 21.9575 30.5995 541 508 1768 ChineseLLaMA-13B 27.7149 22.8383 31.6742 560 490 1768 LLaMA2-7B 27.4321 21.6132 29.8643 528 485 1768 LLaMA-13B 27.0362 20.485 28.6199 506 478 1768 LLaMA-7B 24.6606 20.6606 29.0724 514 436 1768 Table 3: Question Level Result: ”Acc" of human is computed by averaging the human accuracy, and the ’Hit’ of human is determined by the number of questions where the "Acc" of human is above 50%.
Figure 8: Acc&Human Value: Dot line is the human performance, ignore the Human Hit of Human(because it’s 100%)
Based on the evaluation results from ANGO, as shown in Table 3, the majority of participating models performed lower in terms of accuracy compared to existing Chinese benchmarks. This finding indicates that ANGO presents greater challenges for models, thereby enabling more effective identification and performance evaluation.
It is noteworthy that, as shown the ranking details, all English-oriented models performed poorly, and even ChatGPT being outperformed by models with significantly fewer parameters. This outcome can be attributed to ANGO’s emphasis on specific Chinese language features.
Human Value
From Figure 8, there is a positive correlation between Acc and Human Value, indicating that Human Value can indeed reflect the model’s capabilities rather than being arbitrarily fabricated. However, fluctuations are still observed between them within certain ranges. This suggests that Human Value is not a subset or a new expression form of Acc, but rather a distinct indicator. Overall, these quantitative results preliminarily validate that the concept of Human Value captures meaningful signals from a specific evaluative standpoint, affirming that assigning a certain value to model errors can indeed serve as a novel perspective for assessing models.
Based on the results, it can be concluded that models with higher Human Value tend to generate responses that resemble human-like outputs rather than strictly correct answers during conversations. This is because most real humans have limited knowledge in specific domains, and it is precisely this limitation that nurtures specific values and ways of thinking, ultimately shaping different human personalities and traits, which we refer to as "human nature". When the goal is to train AI systems that exhibit "human-like" rather than being "omniscient", incorporating Human Value as a guiding principle may be beneficial.
6 Related Work
The field of English language assessment began its exploration of multiple-choice question benchmarks with the pioneering work of MMLU[6]. MMLU constructed a comprehensive dataset comprising 57 subjects with varying levels of difficulty. This dataset was created by collecting multiple-choice questions in the form of question-answer pairs. Another significant contribution to the evaluation of model truthfulness in question-answering tasks was introduced by TruthfulQA[7]. TruthfulQA transformed traditional question-answering tasks into a multiple-choice format, enabling the assessment of model truthfulness. BIG-Bench[8] took the benchmarking of models to a new level by expanding the classification dimension from two-digit to three-digit levels. It also maintained the dataset dynamically, ensuring its relevance and adaptability.
In the Chinese language domain, CMMLU[11] and MMCU[31] constructed benchmarks inspired by MMLU. These benchmarks were divided based on examination subjects and sources, providing valuable resources for evaluating Chinese language models. AGI-Eval[9] approached benchmarking from a different perspective and optimized it for Chinese characteristics. It created a human-centric benchmark by evaluating model performance based on real human performance, aiming to capture the nuances of Chinese language understanding. Following closely, the release of C-Eval[10] introduced new possibilities for Chinese benchmarks. C-Eval incorporated sophisticated reasoning problems and made efforts to address the issue of data leaks, providing a more comprehensive evaluation platform.
In comparison to existing Chinese domain datasets, ANGO stands out for several distinct features: (1) ANGO includes more complex problems that require models to combine different abilities (2) ANGO incorporates measurements of problem difficulty based on real human performance, using a continuous quantitative metric instead of discrete judgments. (3) Benefited by carefully designed and optimized sampling strategies, the ANGO benchmark enables rapid iteration and minimizes the potential for malicious hacks and data leaks once the leaderboard is launched.
7 Conclusion
The wide coverage and reliability of the ANGO make it a valuable resource that fosters innovation and progress in NLP. By conducting systematic analysis and application of ANGO, researchers can gain a profound understanding of the performance and capabilities of models, thereby providing guidance for model design and improvement. And the unique feature of the ANGO enable developers to better evaluate and enhance their own models. We encourage researchers and developers to actively explore and utilize the ANGO, thereby driving further development and innovation in the field of LLM.
Appendix A Keypoint Chinese-English Mapping
A.1 Expression&Comprehension
Level-1 Chinese Level-1 English Level-2 Chinese Level-2 English Level-3 Chinese Level-3 English Level-4 Chinese Level-4 English
言语理解与表达 Expression&Comprehension
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 关联词-转折 Contrast
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 关联词-因果 Cause And Effect
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 关联词-对策 Countermeasure
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 关联词-并列 Enumeration
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 主题词 Theme Words
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 程度词 Degree Words
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 行文脉络-总分 General To Specific
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 行文脉络-分总 Specific To General
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 行文脉络-分总分 Specific To General To Specific
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 中心理解题 Central Comprehension Questions 特殊问法 Special Question Forms
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 细节判断题 Detail Judgement Questions
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 词句理解题 Word And Sentence Comprehension Questions
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 词句理解题 Word And Sentence Comprehension Questions 实词 Content Words
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 词句理解题 Word And Sentence Comprehension Questions 代词 Pronouns
言语理解与表达 Expression&Comprehension 阅读理解 Reading Comprehension 标题填入题 Title Filling Questions
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 首句特征 First Sentence Features
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 非首句特征 First Sentence Features
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 确定捆绑 Definite Binding
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 确定顺序 Definite Order
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句排序题 Statement Ordering Questions 尾句特征 Ending Sentence Features
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句填空题 Statement Filling Questions
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句填空题 Statement Filling Questions 开头 Beginning
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句填空题 Statement Filling Questions 中间 Middle
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 语句填空题 Statement Filling Questions 结尾 End
言语理解与表达 Expression&Comprehension 语句表达 Statement Expression 接语选择题 Next Sentence Selection Questions
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 实词填空 Content Word Filling
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 成语填空 Idiom Filling
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 混搭填空 Mixed Filling
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 词的辨析 Word Discrimination
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 词的辨析 Word Discrimination 词的辨析-词义侧重 Meaning Emphasis
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 词的辨析 Word Discrimination 词的辨析-固定搭配 Collocations
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 词的辨析 Word Discrimination 词的辨析-感情色彩 Connotations
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 词的辨析 Word Discrimination 词的辨析-程度轻重 Degree
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 语境分析 Context Analysis
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 语境分析 Context Analysis 关联关系-转折关系 Contrast
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 语境分析 Context Analysis 关联关系-因果关系 Cause And Effect
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 语境分析 Context Analysis 关联关系-并列关系 Enumeration
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 语境分析 Context Analysis 对应关系-解释类对应 Explanatory
言语理解与表达 Expression&Comprehension 逻辑填空 Logical Filling 语境分析 Context Analysis 对应关系-重点词句对应 Key Words And Phrases
A.2 Quantitative Relations
Level-1 Chinese Level-1 English Level-2 Chinese Level-2 English Level-3 Chinese Level-3 English Level-4 Chinese Level-4 English
数量关系 Quantitative Relations
数量关系 Quantitative Relations 数学运算 Mathematical Calculations
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 工程问题 Engineering Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 工程问题 Engineering Problems 给完工时间型 Give Completion Time
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 工程问题 Engineering Problems 给效率比例型 Give Efficiency Ratio
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 工程问题 Engineering Problems 给具体单位型 Give Specific Unit
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 工程问题 Engineering Problems 工程问题-其他 Other
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 最值问题 Extreme Value Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 最值问题 Extreme Value Problems 非典型最值问题 Atypical Extreme Value Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 最值问题 Extreme Value Problems 构造数列 Constructing A Sequence
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 最值问题 Extreme Value Problems 最不利构造 Most Unfavorable Construction
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 最值问题 Extreme Value Problems 多集合反向构造 Reverse Construction Of Multiple Sets
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 年龄问题 Age Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 和差倍比问题 Ratio And Difference Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 周期问题 Cyclic Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 周期问题 Cyclic Problems 周期相遇问题 Cyclic Encounter Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 周期问题 Cyclic Problems 周期余数问题 Cyclic Remainder Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 周期问题 Cyclic Problems 周期问题-其他 Other
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 数列问题 Sequence Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems 火车过桥 Train Crossing Bridge
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems 平均速度 Average Speed
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems 普通行程 Regular Journey
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems 相遇追及 Meet And Chase
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems 流水行船 Stream Flow Ship
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 行程问题 Journey Problems 行程问题-其他 Other
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 几何问题 Geometry Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 几何问题 Geometry Problems 平面几何 Plane Geometry
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 几何问题 Geometry Problems 立体几何 Solid Geometry
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 容斥原理问题 Exclusion Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 容斥原理问题 Exclusion Problems 两集合 Two Sets
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 容斥原理问题 Exclusion Problems 三集合 Three Sets
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 基础排列组合 Basic Permutation And Combination
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 相邻问题 Adjacent Issues
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 不相邻问题 Adjacent Issues
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 同素分堆问题 Issues With Same Factors
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 环形排列问题 Circular Arrangement Issues
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 错位排列 Dislocation Arrangement
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 排列组合问题 Permutation And Combination Problems 排列组合问题-其他 Other
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 概率问题 Probability Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 概率问题 Probability Problems 给情况求概率 Given Situation Determine Probability
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 概率问题 Probability Problems 给概率求概率 Given Probability Determine Probability
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 概率问题 Probability Problems 概率问题-其他 Other
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 经济利润问题 Economics And Profit Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 不定方程问题 Indeterminate Equation Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 不定方程问题 Indeterminate Equation Problems 普通不定方程 Ordinary Indeterminate Equations
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 不定方程问题 Indeterminate Equation Problems 不定方程组 Indeterminate Equation Groups
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 统筹规划问题 Coordination Planning Problems
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 数学运算-其他 Other
数量关系 Quantitative Relations 数学运算 Mathematical Calculations 公倍数与公约数问题 Lcm And Gcd Problems
A.3 Judgment&Inference
Level-1 Chinese Level-1 English Level-2 Chinese Level-2 English Level-3 Chinese Level-3 English Level-4 Chinese Level-4 English
判断推理 Judgment&Inference
判断推理 Judgment&Inference 定义判断 Definitions And Judgements
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 主客体 Subject And Object
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 大前提 Major Premise
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 方式目的 Means And Purposes
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 原因结果 Causes And Results
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 单定义-其他句式 Other Patterns
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 故事类 Story
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 单定义 Single Definitions 拆词 Word Parsing
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 多定义 Multiple Definitions
判断推理 Judgment&Inference 定义判断 Definitions And Judgements 多定义 Multiple Definitions 常规问法 Conventional Question Forms
判断推理 Judgment&Inference 逻辑判断 Logical Judgements
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 加强题型 Strengthening Question Types
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 加强题型 Strengthening Question Types 搭桥 Building A Bridge
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 加强题型 Strengthening Question Types 必要条件 Necessary Conditions
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 加强题型 Strengthening Question Types 补充论据 Supplementing Evidence
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 加强题型 Strengthening Question Types 加强选非题 Strengthening Wrong Choice Questions
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 加强题型 Strengthening Question Types 加强-其他 Other
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 削弱论点 Weakening Viewpoints
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 拆桥 Dismantling The Bridge
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 他因削弱 Other Factors Weakening
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 削弱选非题 Weakening Wrong Choice Questions
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 削弱论据 Weakening Evidence
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 因果倒置 Reversing Cause And Effect
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 削弱题型 Weakening Question Types 削弱-其他 Other
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 翻译推理 Translation Reasoning
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 翻译推理 Translation Reasoning 常规翻译 Conventional Translation
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 翻译推理 Translation Reasoning 集合推理 Set Reasoning
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 翻译推理 Translation Reasoning 推理形式 Reasoning Patterns
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 翻译推理 Translation Reasoning 翻译推理-其他 Other
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 组合排列-材料 Materials
判断推理 Judgment&Inference 逻辑判断 Logical Judgements 原因解释 Reason Explanations
判断推理 Judgment&Inference 类比推理 Analogical Reasoning
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 语义关系 Semantic Relationships
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 语义关系 Semantic Relationships 语义关系-近义关系 Synonyms
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 语义关系 Semantic Relationships 语义关系-反义关系 Antonyms
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 语义关系 Semantic Relationships 语义-其他 Other
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 逻辑关系 Logical Relationships
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 逻辑关系 Logical Relationships 逻辑关系-全同关系 Identical
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 逻辑关系 Logical Relationships 逻辑关系-并列关系 Parallel
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 逻辑关系 Logical Relationships 逻辑关系-交叉关系 Intersection
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 逻辑关系 Logical Relationships 逻辑关系-包容关系 Inclusion
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 逻辑关系 Logical Relationships 逻辑关系-对应关系 Correspondence
判断推理 Judgment&Inference 类比推理 Analogical Reasoning 拆分思维 Split Thinking
A.4 Material Analysis
Level-1 Chinese Level-1 English Level-2 Chinese Level-2 English Level-3 Chinese Level-3 English
资料分析 Material Analysis
资料分析 Material Analysis 文字资料 Text Materials
资料分析 Material Analysis 综合资料 Comprehensive Materials
资料分析 Material Analysis 简单计算 Simple Calculations
资料分析 Material Analysis 简单计算 Simple Calculations 直接找数 Directly Finding Numbers
资料分析 Material Analysis 简单计算 Simple Calculations 简单加减计算 Simple Addition And Subtraction Calculations
资料分析 Material Analysis 简单计算 Simple Calculations 排序类 Ordering
资料分析 Material Analysis 基期与现期 Base And Present Periods
资料分析 Material Analysis 基期与现期 Base And Present Periods 基期计算 Base Period Calculations
资料分析 Material Analysis 基期与现期 Base And Present Periods 现期计算 Present Period Calculations
资料分析 Material Analysis 基期与现期 Base And Present Periods 基期比较 Base Period Comparisons
资料分析 Material Analysis 基期与现期 Base And Present Periods 间隔基期 Interval Base Periods
资料分析 Material Analysis 基期与现期 Base And Present Periods 基期和差 Base Period Differences
资料分析 Material Analysis 基期与现期 Base And Present Periods 现期追赶 Present Period Catching Up
资料分析 Material Analysis 增长率 Growth Rates
资料分析 Material Analysis 增长率 Growth Rates 一般增长率 General Growth Rate
资料分析 Material Analysis 增长率 Growth Rates 混合增长率 Compound Growth Rate
资料分析 Material Analysis 增长率 Growth Rates 间隔增长率 Interval Growth Rate
资料分析 Material Analysis 增长率 Growth Rates 年均增长率 Annual Growth Rate
资料分析 Material Analysis 增长量 Growth Amounts
资料分析 Material Analysis 增长量 Growth Amounts 增长量计算 Growth Amount Calculations
资料分析 Material Analysis 增长量 Growth Amounts 增长量比较 Growth Amount Comparisons
资料分析 Material Analysis 增长量 Growth Amounts 间隔增长量 Interval Growth Amounts
资料分析 Material Analysis 增长量 Growth Amounts 年均增长量 Annual Growth Amounts
资料分析 Material Analysis 比重问题 Percentage Problems
资料分析 Material Analysis 比重问题 Percentage Problems 现期比重 Present Period Percentage
资料分析 Material Analysis 比重问题 Percentage Problems 基期比重 Base Period Percentage
资料分析 Material Analysis 比重问题 Percentage Problems 两期比重 Two Period Percentages
资料分析 Material Analysis 比重问题 Percentage Problems 混合比重 Mixed Percentages
资料分析 Material Analysis 平均数问题 Average Problems
资料分析 Material Analysis 平均数问题 Average Problems 基期平均数 Base Period Average
资料分析 Material Analysis 平均数问题 Average Problems 现期平均数 Present Period Average
资料分析 Material Analysis 平均数问题 Average Problems 平均数的增长率 Growth Rate Of Averages
资料分析 Material Analysis 平均数问题 Average Problems 平均数的增长量 Growth Amount Of Averages
资料分析 Material Analysis 平均数问题 Average Problems 两期平均数比较 Comparing Two Period Averages
资料分析 Material Analysis 倍数与比值相关 Multiples And Ratios
资料分析 Material Analysis 倍数与比值相关 Multiples And Ratios 基期倍数 Base Period Multiples
资料分析 Material Analysis 倍数与比值相关 Multiples And Ratios 现期倍数 Present Period Multiples
资料分析 Material Analysis 倍数与比值相关 Multiples And Ratios 比值计算 Ratio Calculations
资料分析 Material Analysis 倍数与比值相关 Multiples And Ratios 比值比较 Ratio Comparisons
资料分析 Material Analysis 综合分析 Comprehensive Analysis
A.5 General Knowledge
Level-1 Chinese Level-1 English Level-2 Chinese Level-2 English Level-3 Chinese Level-3 English
常识判断 General Knowledge
常识判断 General Knowledge 政治常识 Political Knowledge
常识判断 General Knowledge 政治常识 Political Knowledge 时政 Current Affairs
常识判断 General Knowledge 政治常识 Political Knowledge 中国特色社会主义建设 Socialist Construction With Chinese Characteristics
常识判断 General Knowledge 经济常识 Economic Knowledge
常识判断 General Knowledge 经济常识 Economic Knowledge 宏观经济与调控政策 Macroeconomics And Regulation
常识判断 General Knowledge 科技常识 Scientific Knowledge
常识判断 General Knowledge 科技常识 Scientific Knowledge 物理常识 Physics Basics
常识判断 General Knowledge 科技常识 Scientific Knowledge 化学常识 Chemistry Basics
常识判断 General Knowledge 科技常识 Scientific Knowledge 生物常识 Biology Basics
常识判断 General Knowledge 科技常识 Scientific Knowledge 科技理论与成就 Science And Technology Theories And Achievements
常识判断 General Knowledge 科技常识 Scientific Knowledge 生活常识 Common Knowledge
常识判断 General Knowledge 人文常识 Humanities Knowledge
常识判断 General Knowledge 人文常识 Humanities Knowledge 中国历史 Chinese History
常识判断 General Knowledge 人文常识 Humanities Knowledge 世界历史 World History
常识判断 General Knowledge 人文常识 Humanities Knowledge 文学常识 Literature Basics
常识判断 General Knowledge 人文常识 Humanities Knowledge 文化常识 Cultural Basics
常识判断 General Knowledge 地理国情 Geography And National Conditions
常识判断 General Knowledge 地理国情 Geography And National Conditions 自然常识 Natural Science Basics
常识判断 General Knowledge 地理国情 Geography And National Conditions 国情社情 National Conditions
常识判断 General Knowledge 法律常识 Legal Knowledge
常识判断 General Knowledge 法律常识 Legal Knowledge 宪法 Constitution
常识判断 General Knowledge 法律常识 Legal Knowledge 行政法 Administrative Law
常识判断 General Knowledge 法律常识 Legal Knowledge 民法 Civil Law
常识判断 General Knowledge 法律常识 Legal Knowledge 刑法 Criminal Law
常识判断 General Knowledge 法律常识 Legal Knowledge 劳动法 Labor Law
常识判断 General Knowledge 法律常识 Legal Knowledge 其他法律法规 Other Laws And Regulations
常识判断 General Knowledge 法律常识 Legal Knowledge 民事诉讼法 Civil Procedure Law
常识判断 General Knowledge 法律常识 Legal Knowledge 经济法 Economic Law